OpenStudy (anonymous):

how do i solve 5-4/5x=-11

OpenStudy (anonymous):

Subtract 5 from both sides, then multiply both sides by (-5/4) \[5-4/5=-11 so -4/5x=-16 becomes x=-16(-5/4) so x=20\]

OpenStudy (anonymous):

ty so much tdredden i had most of it right except i forgot the negative and switching it ty again